# Shaping non-reciprocal caustic spin-wave beams

### v3

@@ Acknowledgements.
This work has benefited from a government grant operated by the French National Research Agency as part of the France 2030 program, reference ANR-22-EXSP-0004 (SWING), as well as the ANR project MagFunc/ANR-20-CE91-0005, and the Transatlantic Research Partnership, a program of FACE Foundation and the French Embassy. We also acknowledge financial support by the Interdisciplinary Thematic Institute QMat, as part of the ITI 2021-2028 Program of the University of Strasbourg, CNRS and Inserm, IdEx Unistra (ANR 10 IDEX 0002), SFRI STRAT’US Project (ANR 20 SFRI 0012), and ANR-17-EURE-0024 under the framework of the French Investments for the Future Program, as well as the High Performance Computing Center of the University of Strasbourg for supporting this work by providing scientific support and access to computing resources. Part of the computing resources were funded by the Equipex Equip@Meso project (Programme Investissements d’Avenir) and the CPER Alsacalcul/Big Data. Research at the University of Delaware was supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Sciences and Engineering under Award DE-SC0020308. The authors acknowledge the use of facilities and instrumentation supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. We thank Anish Rai for characterizing the YIG films by FMR.References
- Kravtsov and Orlov [1993] Y. A. Kravtsov and Y. I. Orlov, Caustics, Catastrophes and Wave Fields (Springer Berlin Heidelberg, 1993).
- Davydov and Troitsky [2023] D. Davydov and S. Troitsky, Testing universal dark-matter caustic rings with galactic rotation curves, Physics Letters B 839, 137798 (2023).
- White and Vogelsberger [2009] S. D. M. White and M. Vogelsberger, Dark matter caustics, Monthly Notices of the Royal Astronomical Society 392, 281 (2009).
- Every [1986] A. G. Every, Formation of phonon-focusing caustics in crystals, Physical Review B 34, 2852 (1986).
- Maris [1971] H. J. Maris, Enhancement of heat pulses in crystals due to elastic anisotropy, The Journal of the Acoustical Society of America 50, 812–818 (1971).
- Taylor et al. [1969] B. Taylor, H. J. Maris, and C. Elbaum, Phonon focusing in solids, Physical Review Letters 23, 416–419 (1969).
- Shi et al. [2015] X. Shi, X. Lin, F. Gao, H. Xu, Z. Yang, and B. Zhang, Caustic graphene plasmons with Kelvin angle, Physical Review B 92, 081404 (2015).
- Epstein and Arie [2014] I. Epstein and A. Arie, Arbitrary bending plasmonic light waves, Physical Review Letters 112, 023903 (2014).
- Spector et al. [1990] J. Spector, H. L. Stormer, K. W. Baldwin, L. N. Pfeiffer, and K. W. West, Electron focusing in two-dimensional systems by means of an electrostatic lens, Applied Physics Letters 56, 1290–1292 (1990).
- Cheianov et al. [2007] V. V. Cheianov, V. Fal’ko, and B. L. Altshuler, The focusing of electron flow and a veselago lens in graphene p-n junctions, Science 315, 1252–1255 (2007).
- Cserti et al. [2007] J. Cserti, A. Pályi, and C. Péterfalvi, Caustics due to a negative refractive index in circular graphene junctions, Physical Review Letters 99, 10.1103/physrevlett.99.246801 (2007).
@@
- Büttner et al. [2000] O. Büttner, M. Bauer, S. O. Demokritov, B. Hillebrands, Y. S. Kivshar, V. Grimalsky, Y. Rapoport, and A. N. Slavin, Linear and nonlinear diffraction of dipolar spin waves in yttrium iron garnet films observed by space- and time-resolved brillouin light scattering, Physical Review B 61, 11576–11587 (2000).
- Veerakumar and Camley [2006] V. Veerakumar and R. E. Camley, Magnon focusing in thin ferromagnetic films, Physical Review B 74 (2006).
@@
- Demidov et al. [2009] V. E. Demidov, S. O. Demokritov, D. Birt, B. O’Gorman, M. Tsoi, and X. Li, Radiation of spin waves from the open end of a microscopic magnetic-film waveguide, Physical Review B 80 (2009).
- Schneider et al. [2010] T. Schneider, A. A. Serga, A. V. Chumak, C. W. Sandweg, S. Trudel, S. Wolff, M. P. Kostylev, V. S. Tiberkevich, A. N. Slavin, and B. Hillebrands, Nondiffractive subwavelength wave beams in a medium with externally controlled anisotropy, Physical Review Letters 104 (2010).
- Kostylev et al. [2011] M. P. Kostylev, A. A. Serga, and B. Hillebrands, Radiation of caustic beams from a collapsing bullet, Physical Review Letters 106, 134101 (2011).
- Sebastian et al. [2013] T. Sebastian, T. Brächer, P. Pirro, A. A. Serga, B. Hillebrands, T. Kubota, H. Naganuma, M. Oogane, and Y. Ando, Nonlinear emission of spin-wave caustics from an edge mode of a microstructured Co2Mn0.6Fe0.4Si waveguide, Physical Review Letters 110, 067201 (2013).
- Gieniusz et al. [2013] R. Gieniusz, H. Ulrichs, V. D. Bessonov, U. Guzowska, A. I. Stognii, and A. Maziewski, Single antidot as a passive way to create caustic spin-wave beams in yttrium iron garnet films, Applied Physics Letters 102, 102409 (2013).
- Madami et al. [2018] M. Madami, Y. Khivintsev, G. Gubbiotti, G. Dudko, A. Kozhevnikov, V. Sakharov, A. Stal’makhov, A. Khitun, and Y. Filimonov, Nonreciprocity of backward volume spin wave beams excited by the curved focusing transducer, Applied Physics Letters 113 (2018).
- Shiota et al. [2020] Y. Shiota, S. Funada, R. Hisatomi, T. Moriyama, and T. Ono, Imaging of caustic-like spin wave beams using optical heterodyne detection, Applied Physics Letters 116, 192411 (2020).
- Muralidhar et al. [2021] S. Muralidhar, R. Khymyn, A. A. Awad, A. Alemán, D. Hanstorp, and J. Åkerman, Femtosecond laser pulse driven caustic spin wave beams, Physical Review Letters 126, 037204 (2021).
- Gallardo et al. [2021] R. A. Gallardo, P. Alvarado-Seguel, A. Kákay, J. Lindner, and P. Landeros, Spin-wave focusing induced by dipole-dipole interaction in synthetic antiferromagnets, Physical Review B 104 (2021).
- Wartelle et al. [2023] A. Wartelle, F. Vilsmeier, T. Taniguchi, and C. H. Back, Caustic spin wave beams in soft thin films: Properties and classification, Phys. Rev. B 107, 144431 (2023).
- Kim et al. [2016] J.-V. Kim, R. L. Stamps, and R. E. Camley, Spin wave power flow and caustics in ultrathin ferromagnets with the dzyaloshinskii-moriya interaction, Physical Review Letters 117, 10.1103/physrevlett.117.197204 (2016).
- Papp et al. [2021a] A. Papp, G. Csaba, and W. Porod, Characterization of nonlinear spin-wave interference by reservoir-computing metrics, Applied Physics Letters 119, 112403 (2021a).
- Lee and Mochizuki [2022] M.-K. Lee and M. Mochizuki, Reservoir computing with spin waves in a skyrmion crystal, Physical Review Applied 18 (2022).
- Lee et al. [2023] O. Lee, T. Wei, K. D. Stenning, J. C. Gartside, D. Prestwood, S. Seki, A. Aqeel, K. Karube, N. Kanazawa, Y. Taguchi, C. Back, Y. Tokura, W. R. Branford, and H. Kurebayashi, Task-adaptive physical reservoir computing, Nature Materials (2023).
- Khitun [2013] A. Khitun, Magnonic holographic devices for special type data processing, Journal of Applied Physics 113, 164503 (2013).
- Khitun [2015] A. Khitun, Parallel database search and prime factorization with magnonic holographic memory devices, Journal of Applied Physics 118, 243905 (2015).
- Papp et al. [2021b] Á. Papp, W. Porod, and G. Csaba, Nanoscale neural network using non-linear spin-wave interference, Nature Communications 12 (2021b).
- Christensen et al. [2022] D. V. Christensen, R. Dittmann, B. Linares-Barranco, A. Sebastian, M. Le Gallo, A. Redaelli, S. Slesazeck, T. Mikolajick, S. Spiga, S. Menzel, I. Valov, G. Milano, C. Ricciardi, S.-J. Liang, F. Miao, M. Lanza, T. J. Quill, S. T. Keene, A. Salleo, J. Grollier, D. Marković, A. Mizrahi, P. Yao, J. J. Yang, G. Indiveri, J. P. Strachan, S. Datta, E. Vianello, A. Valentian, J. Feldmann, X. Li, W. H. P. Pernice, H. Bhaskaran, S. Furber, E. Neftci, F. Scherr, W. Maass, S. Ramaswamy, J. Tapson, P. Panda, Y. Kim, G. Tanaka, S. Thorpe, C. Bartolozzi, T. A. Cleland, C. Posch, S. Liu, G. Panuccio, M. Mahmud, A. N. Mazumder, M. Hosseini, T. Mohsenin, E. Donati, S. Tolu, R. Galeazzi, M. E. Christensen, S. Holm, D. Ielmini, and N. Pryds, 2022 roadmap on neuromorphic computing and engineering, Neuromorphic Computing and Engineering 2, 022501 (2022).
- Bertelli et al. [2020] I. Bertelli, J. J. Carmiggelt, T. Yu, B. G. Simon, C. C. Pothoven, G. E. W. Bauer, Y. M. Blanter, J. Aarts, and T. van der Sar, Magnetic resonance imaging of spin-wave transport and interference in a magnetic insulator, Science Advances 6, eabd3556 (2020).
- Makartsou et al. [2024] U. Makartsou, M. Golebiewski, U. Guzowska, A. Stognij, R. Gieniusz, and M. Krawczyk, Spin-wave self-imaging: Experimental and numerical demonstration of caustic and talbot-like diffraction patterns, Applied Physics Letters 124, 10.1063/5.0195099 (2024).
- Vlaminck et al. [2023] V. Vlaminck, L. Temdie, V. Castel, M. B. Jungfleisch, D. Stoeffler, Y. Henry, and M. Bailleul, Spin wave diffraction model for perpendicularly magnetized films, Journal of Applied Physics 133, 053903 (2023).
- Temdie et al. [2024] L. Temdie, V. Castel, M. Jungfleisch, R. Bernard, H. Majjad, D. Stoeffler, Y. Henry, M. Bailleul, and V. Vlaminck, Probing spin wave diffraction patterns of curved antennas, Phys. Rev. Appl. 21, 014032 (2024).
- B A Kalinikos and A N Slavin [1986] B A Kalinikos and A N Slavin, Theory of dipole-exchange spin wave spectrum for ferromagnetic films with mixed exchange boundary conditions, Journal of Physics C: Solid State Physics 19, 7013 (1986).
- Kalinikos [1981] B. A. Kalinikos, Spectrum and linear excitation of spin waves in ferromagnetic films, Soviet Physics Journal 24, 718 (1981).
- Kalinikos and Slavin [1986] B. A. Kalinikos and A. N. Slavin, Theory of dipole-exchange spin wave spectrum for ferromagnetic films with mixed exchange boundary conditions, Journal of Physics C: Solid State Physics 19, 7013 (1986).
- SM/ [a] (a), see section I-NFD MODEL FOR IN-PLANE MAGNETIZED THIN FILM of the Supplemental Material at URL-will-be-inserted-by-publisher for the data of the experiments.
- Yu and Bauer [2021] T. Yu and G. E. W. Bauer, Chiral coupling to magnetodipolar radiation, in Chirality, Magnetism and Magnetoelectricity: Separate Phenomena and Joint Effects in Metamaterial Structures, edited by E. Kamenetskii (Springer International Publishing, Cham, 2021) pp. 1–23.
@@
- Devolder [2023] T. Devolder, Propagating-spin-wave spectroscopy using inductive antennas: Conditions for unidirectional energy flow, Physical Review Applied 20, 10.1103/physrevapplied.20.054057 (2023).
- Temdie et al. [2023a] L. Temdie, V. Castel, C. Dubs, G. Pradhan, J. Solano, H. Majjad, R. Bernard, Y. Henry, M. Bailleul, and V. Vlaminck, High wave vector non-reciprocal spin wave beams, AIP Advances 13, 025207 (2023a).
- Temdie et al. [2023b] L. Temdie, V. Castel, T. Reimann, M. Lindner, C. Dubs, G. Pradhan, J. Solano, R. Bernard, H. Majjad, Y. Henry, M. Bailleul, and V. Vlaminck, Chiral excitation of exchange spin waves using gold nanowire grating, Magnetochemistry 9, 199 (2023b).
- SM/ [b] (b), see section II-CHIRAL EXCITATION OF CAUSTICS BEAM of the Supplemental Material at URL-will-be-inserted-by-publisher for the data of the experiments.
@@
- Gurevich and Melkov [1996] A. G. Gurevich and G. A. Melkov, Magnetization Oscillations and Waves (CRC Press, Boca Raton, FL, 1996).
- SM/ [c] (c), see section VI-BRILLOUIN LIGHT SCATTERING SPECTROSCOPY of the Supplemental Material at URL-will-be-inserted-by-publisher for the data of the experiments.
- SM/ [d] (d), see section III-MICROWAVE FIELD DISTRIBUTION of the Supplemental Material at URL-will-be-inserted-by-publisher for the data of the experiments.
- Sebastian et al. [2015] T. Sebastian, K. Schultheiss, B. Obry, B. Hillebrands, and H. Schultheiss, Micro-focused brillouin light scattering: imaging spin waves at the nanoscale, Frontiers in Physics 3 (2015).
- Chumak et al. [2015] A. V. Chumak, V. I. Vasyuchka, A. A. Serga, and B. Hillibrands, Nat. Phys. 11, 453 (2015).
- Csaba et al. [2017] G. Csaba, Á. Papp, and W. Porod, Perspectives of using spin waves for computing and signal processing, Physics Letters A 381, 1471 (2017).
- Pirro et al. [2021] P. Pirro, V. I. Vasyuchka, A. A. Serga, and B. Hillebrands, Advances in coherent magnonics, Nature Reviews Materials 6, 1114 (2021).
- [52] M. Benjamin Jungfleisch, Inelastic scattering of light by spin waves, in Optomagnonic Structures, Chap. 4, pp. 175–211.
- Vansteenkiste et al. [2014] A. Vansteenkiste, J. Leliaert, M. Dvornik, M. Helsen, F. Garcia-Sanchez, and B. Van Waeyenberge, The design and verification of MuMax3, AIP Advances 4, 107133 (2014).
- Kostylev [2013] M. Kostylev, Non-reciprocity of dipole-exchange spin waves in thin ferromagnetic films, Journal of Applied Physics 113 (2013).
- Haidar et al. [2014] M. Haidar, M. Bailleul, M. Kostylev, and Y. Lao, Nonreciprocal oersted field contribution to the current-induced frequency shift of magnetostatic surface waves, Physical Review B 89 (2014).

### v4

@@ Acknowledgements.
This work has benefited from a government grant operated by the French National Research Agency as part of the France 2030 program, Reference No. ANR-22-EXSP-0004 (SWING), as well as the ANR project MagFunc/ANR-20-CE91-0005, and the Transatlantic Research Partnership, a program of FACE Foundation and the French Embassy. We also acknowledge financial support by the Interdisciplinary Thematic Institute QMat, as part of the ITI 2021-2028 Program of the University of Strasbourg, CNRS and Inserm, IdEx Unistra (ANR 10 IDEX 0002), SFRI STRAT’US Project (ANR 20 SFRI 0012), and ANR-17-EURE-0024 under the framework of the French Investments for the Future Program, as well as the High Performance Computing Center of the University of Strasbourg for supporting this work by providing scientific support and access to computing resources. Part of the computing resources were funded by the Equipex Equip@Meso project (Programme Investissements d’Avenir) and the CPER Alsacalcul/Big Data. Research at the University of Delaware was supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Sciences and Engineering under Award No. DE-SC0020308. The authors acknowledge the use of facilities and instrumentation supported by NSF through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. We thank Anish Rai for characterizing the YIG films by FMR.References
- [1] Note: See Supplemental Material at http://link.aps.org/supplemental/DOI for: Sec. I, NFD model for in-plane magnetized thin film ; Sec. II, chiral excitation of caustics beam; Sec. III, microwave field distribution; Sec. VI, Brillouin light scattering spectroscopy. Cited by: Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams.
- [2] (2020-11) Magnetic resonance imaging of spin-wave transport and interference in a magnetic insulator. Sci. Adv. 6 (46), pp. eabd3556. Cited by: Shaping non-reciprocal caustic spin wave beams.
@@
- [3] (2000-05) Linear and nonlinear diffraction of dipolar spin waves in yttrium iron garnet films observed by space- and time-resolved brillouin light scattering. Phys. Rev. B 61, pp. 11576–11587. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [4] (2007-03) The focusing of electron flow and a veselago lens in graphene p-n junctions. Science 315 (5816), pp. 1252–1255. External Links: ISSN 1095-9203, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [5] (2022-05) 2022 roadmap on neuromorphic computing and engineering. Neuromorphic Computing and Engineering 2 (2), pp. 022501. External Links: ISSN 2634-4386 Cited by: Shaping non-reciprocal caustic spin wave beams.
- [6] (2007-12) Caustics due to a negative refractive index in circular graphene junctions. Phys. Rev. Lett. 99, pp. 246801. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [7] (2023) Testing universal dark-matter caustic rings with galactic rotation curves. Physics Letters B 839, pp. 137798. External Links: ISSN 0370-2693, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
@@
- [8] (2009-07) Radiation of spin waves from the open end of a microscopic magnetic-film waveguide. Physical Review B 80 (1), pp. 014429. External Links: ISSN 1550-235X Cited by: Shaping non-reciprocal caustic spin wave beams.
@@
- [9] (2023-11) Propagating-spin-wave spectroscopy using inductive antennas: conditions for unidirectional energy flow. Phys. Rev. Appl. 20, pp. 054057. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [10] (2014-01) Arbitrary bending plasmonic light waves. Physical Review Letters 112 (2), pp. 023903. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [11] (1986-08) Formation of phonon-focusing caustics in crystals. Physical Review B 34 (4), pp. 2852–2862. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [12] (2021-11) Spin-wave focusing induced by dipole-dipole interaction in synthetic antiferromagnets. Phys. Rev. B 104, pp. 174417. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [13] (2013-03) Single antidot as a passive way to create caustic spin-wave beams in yttrium iron garnet films. Applied Physics Letters 102 (10), pp. 102409. External Links: ISSN 0003-6951, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
@@
- [14] (1996-09) Magnetization oscillations and waves. CRC Press, Boca Raton. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [15] (2018-12) Frequency‐division multiplexing in magnonic logic networks based on caustic‐like spin‐wave beams. Phys. Status Solidi Rapid Res. Lett. 12 (12), pp. 1800409. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [16] (1986-12) Theory of dipole-exchange spin wave spectrum for ferromagnetic films with mixed exchange boundary conditions. Journal of Physics C: Solid State Physics 19 (35), pp. 7013–7033. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams.
- [17] (1981-08) Spectrum and linear excitation of spin waves in ferromagnetic films. Soviet Physics Journal 24 (8), pp. 718–731. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [18] (2013-04) Magnonic holographic devices for special type data processing. Journal of Applied Physics 113 (16), pp. 164503. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [19] (2015-12) Parallel database search and prime factorization with magnonic holographic memory devices. Journal of Applied Physics 118 (24), pp. 243905. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [20] (2016-11) Spin wave power flow and caustics in ultrathin ferromagnets with the dzyaloshinskii-moriya interaction. Physical Review Letters 117 (19). External Links: ISSN 1079-7114, Link, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [21] (2011-03) Radiation of caustic beams from a collapsing bullet. Physical Review Letters 106 (13), pp. 134101. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [22] (1993) Caustics, catastrophes and wave fields. Springer Berlin. External Links: ISBN 9783642974915, ISSN 0931-7252, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [23] (2022-07) Reservoir computing with spin waves in a skyrmion crystal. Phys. Rev. Appl. 18, pp. 014074. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [24] (2024-01) Task-adaptive physical reservoir computing. Nature Materials 23 (1), pp. 79–87. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [25] (2018-10) Nonreciprocity of backward volume spin wave beams excited by the curved focusing transducer. Applied Physics Letters 113 (15), pp. 152403. External Links: ISSN 0003-6951, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [26] (2024-05) Spin-wave self-imaging: experimental and numerical demonstration of caustic and talbot-like diffraction patterns. Applied Physics Letters 124 (19), pp. 192406. External Links: ISSN 1077-3118, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [27] (1971-09) Enhancement of heat pulses in crystals due to elastic anisotropy. The Journal of the Acoustical Society of America 50 (3B), pp. 812–818. External Links: ISSN 1520-8524, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [28] (2021-01) Femtosecond laser pulse driven caustic spin wave beams. Physical Review Letters 126 (3), pp. 037204. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [29] (2021-09) Characterization of nonlinear spin-wave interference by reservoir-computing metrics. Applied Physics Letters 119 (11), pp. 112403. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [30] (2021-11) Nanoscale neural network using non-linear spin-wave interference. Nature Commun. 12 (1), pp. 6422. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [31] (2010-05) Nondiffractive subwavelength wave beams in a medium with externally controlled anisotropy. Phys. Rev. Lett. 104, pp. 197203. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [32] (2013-02) Nonlinear emission of spin-wave caustics from an edge mode of a microstructured Co2Mn0.6Fe0.4Si waveguide. Physical Review Letters 110 (6), pp. 067201. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [33] (2015-08) Caustic graphene plasmons with Kelvin angle. Physical Review B 92 (8), pp. 081404. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [34] (2020-05) Imaging of caustic-like spin wave beams using optical heterodyne detection. Applied Physics Letters 116 (19), pp. 192411. External Links: ISSN 0003-6951, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [35] (1990-03) Electron focusing in two-dimensional systems by means of an electrostatic lens. Applied Physics Letters 56 (13), pp. 1290–1292. External Links: ISSN 1077-3118, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [36] (1969-08) Phonon focusing in solids. Physical Review Letters 23 (8), pp. 416–419. External Links: ISSN 0031-9007, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [37] (2023-02) High wave vector non-reciprocal spin wave beams. AIP Advances 13 (2), pp. 025207. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [38] (2024-01) Probing spin wave diffraction patterns of curved antennas. Phys. Rev. Appl. 21, pp. 014032. External Links: Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [39] (2023-08) Chiral excitation of exchange spin waves using gold nanowire grating. Magnetochem. 9 (8), pp. 199. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [40] (2014-10) The design and verification of MuMax3. AIP Advances 4 (10), pp. 107133. External Links: ISSN 2158-3226, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [41] (2006-12) Magnon focusing in thin ferromagnetic films. Phys. Rev. B 74, pp. 214401. External Links: Document, Link Cited by: Shaping non-reciprocal caustic spin wave beams.
- [42] (2023-02) Spin wave diffraction model for perpendicularly magnetized films. Journal of Applied Physics 133 (5), pp. 053903. External Links: ISSN 0021-8979, Document Cited by: Shaping non-reciprocal caustic spin wave beams, Shaping non-reciprocal caustic spin wave beams.
- [43] (2023-04) Caustic spin wave beams in soft thin films: properties and classification. Phys. Rev. B 107, pp. 144431. Cited by: Shaping non-reciprocal caustic spin wave beams.
- [44] (2009-01) Dark matter caustics. Monthly Notices of the Royal Astronomical Society 392 (1), pp. 281–286. External Links: ISSN 0035-8711, Document Cited by: Shaping non-reciprocal caustic spin wave beams.
- [45] (2021) Chiral coupling to magnetodipolar radiation. In Topics in Applied Physics, Topics in Applied Physics, pp. 1–23. Cited by: Shaping non-reciprocal caustic spin wave beams.